OpenStudy (anonymous):

lim_{x rightarrow -infty} {{4 - 7x} over {2 + 3x}}

OpenStudy (anonymous):

\[\lim_{x \rightarrow -\infty} {{4 - 7x} \over {2 + 3x}}\]

OpenStudy (anonymous):

im gonna say lopital's rule.. so -7/3. im sure its wrong though, i cant find anything to factor out.

OpenStudy (anonymous):

factor an x from the top and bottom.

OpenStudy (anonymous):

\[\frac{x(-7+\frac{4}{x})}{x(3+\frac{2}{x})}\]

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

So you should get \(\frac{-7}{3}\) when you cancel the x and take the limit because the two fractions on top and bottom will go to 0.

OpenStudy (anonymous):

Right. Now i can go on. thks